# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ 3 Asymptotic Theory
Under Assumptions 14, there exists some absolute constant $C_{{\rm m}}$ such that when the LASSO tuning parameter (2.2) for the main regression satisfies $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$, as $n\to\infty$, we have
@@
Lemma 1 provides the consistency of the standardized coefficients, which is necessary to establish the asymptotic normality of XDlasso. The rates of tuning parameters are technical devices for proofs. In practice, we recommend cross-validation for the tuning parameter selection.
@@ 3.1 The Auxiliary Regression
While the convergence of Slasso in the main regression is well expected, new theory must be developed for the auxiliary regression (2.10), again by the workhorse Slasso when $w_{j,t}$ is either a unit root or stationary regressor. Importantly, as we maintain an agnostic attitude about the persistence of the variable of interest, which is the dependent variable of the auxiliary regression, the theory must be simultaneously applicable to the two very different types of time series in a unified manner. Without loss of generality, let $\mathcal{M}_{x}=[p_{x}]$ and $\mathcal{M}_{z}=[p]\backslash[p_{x}]$ denote the integer sets indexing the locations of unit root and stationary regressors, respectively.
We first examine the case when $j\in\mathcal{M}_{x}.$ The standardized instrument $\widetilde{\zeta}_{j,t}$ has a different degree of persistence from $X_{-j,t}$ and $Z_{t}$, due to the different orders of integration. In the low dimensional framework, the FCLT in Phillips and Lee, (2016) yields the following rate of convergence in OLS:
@@
$\widetilde{\zeta}_{j,t}=X_{t}^{\top}0_{p_{x}}+Z_{-j,t}^{\top}\varphi_{z}^{(j)*}+\widetilde{\eta}_{j,t}=Z_{-j,t}^{\top}\varphi_{z}^{(j)*}+\widetilde{\eta}_{j,t},$ (3.10)
[…]
In addition, the error term $\widetilde{\eta}_{j,t}$ is close to the stationary time series $(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$ and thus asymptotically uncorrelated to the nonstationary regressors $X_{t}$ due to different persistence. Furthermore, the coefficients $\varphi_{0z}^{(j)*}$ satisfy $\mathbb{E}\left(Z_{-j,t}(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})\right)=0$. Therefore, $\widetilde{\eta}_{j,t}$ is also asymptotically uncorrelated to the the stationary regressors $Z_{-j,t}$, thereby ensuring the consistency of the LASSO estimator $\widehat{\varphi}^{(j)}$ in (2.10).
We impose the following assumption on the coefficient $\varphi_{0z}^{(j)*}$.
@@
$\|\varphi_{0z}^{(j)*}\|_{0}\leq s$ with $\varphi_{0z}^{(j)*}$ defined as (3.8) and $s$ specified in Assumption 4. Moreover, $\|\varphi_{0z}^{(j)*}\|_{1}\leq C_{1}$ for some absolute constant $C_{1}.$
To bound the LASSO estimation errors, we need sparsity of not only the main regression (2.1), but also the auxiliary regression (3.10). This is similar to the high dimensional sparse instrumental variable regression; see Zhu, (2018) and Gold et al., (2020). In Assumption 5, we slightly abuse the sparsity index $s$ to bound the number of nonzero coefficients in the vector $\varphi_{0z}^{(j)*}$. For simplicity, we directly impose this high-level sparsity assumption on $\varphi_{0z}^{(j)*}$, which can be deduced under the commonly used sparsity restriction on the precision matrix $\left(\mathbb{E}\left(Z_{t}Z_{t}^{\top}\right)\right)^{-1}$; see the definition of $s_{j}$ and the conditions in Theorem 2.1 of Zhang and Cheng, (2017, pp. 759). Finally, the upper bound of the $L_{1}$-norm of $\varphi_{0z}^{(j)*}$ controls the variance of the error term in the pseudo-true model (3.10). Parallel conditions naturally hold for $j\in\mathcal{M}_{x}$ as the pseudo-true coefficients are zero in view of the local-to-zero OLS estimate displayed in (3.7).
The following proposition formally lays out the convergence rate of the auxiliary LASSO estimator (2.10).
@@
Suppose that $j\in\mathcal{M}_{z}$ and Assumptions 15 hold. Then, there exists some absolute constant $C_{{\rm a}}>0$ such that when $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, we have
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right)$ (3.11)
@@
$\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t}\widehat{r}_{j,t}\|_{\infty}\leq\dfrac{\mu}{2}\leq\dfrac{C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (3.12)
which is sufficient to bound the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$ in (2.5). Nevertheless, it is still necessary to establish the consistency of $\widehat{\varphi}^{(j)}$. This consistency result not only helps us show the asymptotic normality of the $t$-statistic $t_{j}^{{\rm XD}}$ defined in (2.13), which guarantees the asymptotic size of the test, but is also critical for the convergence rate of the standard error that governs the asymptotic power.
These key theoretical results about asymptotic size and power are displayed in the next section.
3.2 Asymptotic Normality
The consistency of the auxiliary regression built in Proposition 1 guarantees the following decomposition of the statistic:
$\frac{\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}}{\widehat{\omega}_{j}^{{\rm XD}}}={\rm sgn}_{j}\cdot\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}+O_{p}\left(\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\right),$ (3.13)
where ${\rm sgn}_{j}=\frac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is either 1 or $-1$ with probability one as $\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}$ is continuously distributed. The first term of the above expression is a counterpart of $\mathrm{N}_{j}/\omega_{j}$ in the discussion of generic desparsifying argument in (2.5). The second term is the convergence rate of the approximation error of the shrinkage bias analogous to $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}$. Under Assumption 4, the second term of the right-hand side of (3.13) is asymptotically negligible, thereby yielding the following asymptotic normality for any $j\in[p]$.
@@ 3.2 Asymptotic Normality
Suppose Assumptions 1-5 hold. There exist absolute constants $C_{{\rm m}}$ and $C_{{\rm a}}$ such that when $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, respectively, as $n\to\infty$, we have
@@
$\widehat{\omega}_{j}^{{\rm XD}}=\boldsymbol{1}\left\{j\in\mathcal{M}_{x}\right\}\cdot O_{p}\left(1/\sqrt{n^{1+\tau}}\right)+\boldsymbol{1}\left\{j\in\mathcal{M}_{z}\right\}\cdot O_{p}\left(1/\sqrt{n}\right).$
The convergence rates displayed in Theorem 2 are coherent with the results in low dimensional predictive regressions. When $j\in\mathcal{M}_{x}$, the standard error converges faster than the rate $1/n^{\delta_{j}}$ for any $\delta_{j}\in(0,(1+\tau)/2)$. Thus, for the null hypothesis $\mathbb{H}_{0}:\beta_{j}^{*}=0$, the hypothesis testing based on the $t$-statistic $t_{j}^{{\rm XD}}$ is consistent under the alternative with $\beta_{j}^{*}=c/n^{\delta_{j}}$ for some $c>0$ over this class of $\delta_{j}$. The range $\delta_{j}\in(0,(1+\tau)/2)$ includes the important $1/\sqrt{n}$ rate of coefficients for a unit root $X_{j,t}$, under which ${\rm var}(X_{j,t}\beta_{j}^{*})=O(1).$ The $1/\sqrt{n}$ factor thus balances the larger order of unit root regressors $X_{t}$ and the standard $O_{p}(1)$ stochastic order of $y_{t}$ (Phillips,, 2015; Lee et al.,, 2022), and the test by XDlasso is consistent under this class of alternatives. When $j\in\mathcal{M}_{z}$, XDlasso achieves the standard $\sqrt{n}$-consistency.
@@
Unlike the convergence rates of Slasso in Lemma 1, the convergence rate of XDlasso in Theorem 2 is independent of the variable dimension $p$ and the sparsity index $s$. In view of (2.5), in XDlasso the order of the approximation error of the shrinkage bias, $\widehat{{\rm B}}_{j}-{\rm B}_{j}$, depends on $p$ and $s$, but it is dominated by the order of the noise component ${\rm N}_{j}$. The convergence rate of XDlasso therefore follows the order of ${\rm N}_{j}$, which relates only to the sample size $n$.
Remark 8.
Our theoretical results do not cover cointegrated data. Mei and Shi, (2024, Theorem 4) show that in the presence of cointegration, Slasso over-penalizes the coefficients of cointegrated regressors and shrinks them all the way to zero, regardless of their true values. We are unaware of any regularization method that achieves consistent estimation in the $p\gg n$ scheme with cointegrated variables mixed with unit roots and stationary ones. Without consistency, inference about the cointegrated variable’s true parameter is beyond reach at this moment.
Let us summarize the insights gained from the theoretical development. The fundamental principle behind the Dlasso method, as discussed in Zhang and Zhang, (2014), is based on the Frisch-Waugh-Lovell theorem. This theorem purges the influence of other control variables to overcome shrinkage bias. In the context of persistent predictors, the two-stage least squares approach, as developed in Magdalinos and Phillips, (2009) is used to eliminate Stambaugh bias. Each piece is a modern version of a classical idea. To adapt these existing procedures to high dimensional predictive regressions, we must rely on the consistency of Slasso, as proved in Mei and Shi, (2024). This consistency is crucial for both the main regression and the auxiliary regression. Moreover, as mentioned in Remark 5, the construction of the IV and the choice of $\tau$ must be tailored to balance the two types of predictors. This approach is original in that such a restriction is unique to high dimensional models and has not been studied before even in conventional low-dimensional settings.
4 Simulation
In this section, we evaluate the performance of the proposed XDlasso inference procedure by comparing its test size and power with those of Dlasso. Although Dlasso, as well as other existing LASSO-bias-correction procedures, is not designed to handle persistent regressors, we include this comparison to highlight the value added by the IVX transformation. Furthermore, to demonstrate the robustness of the XDlasso approach, we benchmark it using tests based on infeasible estimators where an oracle reveals the locations of nonzero coefficients.

### v3

@@ 3.1 Consistency of Slasso
Under Assumptions 14, there exists an absolute constant $c_{\kappa}$ such that
$\widehat{\kappa}_{D}\geq\frac{c_{\kappa}}{s(\log p)^{4}}\cdot$ (3.7)
w.p.a.1 as $n\to\infty.$ In addition, there exists some absolute constant $C_{{\rm DB}}$ sch that
$4\,\bigg\|\dfrac{1}{n}\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\bigg\|_{\infty}\leq\dfrac{C_{{\rm DB}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}$ (3.8)
w.p.a.1 as $n\to\infty$, where $r$ is defined in Assumption 2.
Since LURs share similar asymptotic behavior with unit roots, the orders of the RE (Eq. (3.7)) and DB (Eq. (3.8)) are the same as those in Proposition 3 of MS24. Nevertheless, the technical proofs for LURs are challenging.
Remark 10.
For illustration, in this remark we suppose that all regressors are LURs. In low dimensions, the Gram matrix of LURs, after scaled by $n^{-1}$, converges in distribution to a non-degenerate stochastic integral
$n^{-1}\widehat{\Sigma}\stackrel{{\scriptstyle{\rm d}}}{{\to}}\int_{0}^{1}\mathcal{\mathcal{U}}_{{\rm{\bf C}}^{*}}(r)\mathcal{\mathcal{\mathcal{U}}}_{{\rm{\bf C}}^{*}}(r)^{\top}{\rm d}r,$ (3.9)
where $\mathcal{\mathcal{\mathcal{U}}}_{{\rm{\bf{\rm{\bf C}}^{*}}}}(t):=\int_{0}^{t}{\rm e}^{{\rm{\bf{\rm{\bf C}}^{*}}}(t-r)}{\rm d}\mathcal{B}(r)$ is a vector of Ornstein–Uhlenbeck processes, with ${\rm{\bf C}}^{*}:={\rm diag}(c_{1}^{*},c_{2}^{*},\cdots,c_{p_{x}}^{*})$ storing the constants in the AR(1) coefficients of LURs in (2.3), and $\mathcal{B}(r)$ being a multivariate Brownian motion. The diagonal entries of the stochastic integral on the right-hand side of (3.9) are nonnegative and continuously distributed, with a non-trivial probability in a neighborhood of zero. Consequently, the minimum diagonal entry of $n^{-1}\widehat{\Sigma}$ diminishes to zero as the dimension of LUR regressors passes to infinity. Eq. (3.7) establishes a lower bound of RE that shrinks to zero in a sufficiently slow speed, thereby still ensuring the consistency of Slasso. Under LUR, the linear coefficients in $x_{j,t}=\sum_{\ell\geq 0}(1+c_{j}^{*}/n)^{\ell}e_{j,t-\ell}$ depends on $n$ and $\ell$, which is much more complicated to deal with than the special case $c_{j}^{*}=0$ in MS24, where all the linear coefficients become $1$ and the representation of $x_{j,t}$ becomes a simple partial sum of stationary components. Interested readers may refer to Lemma A.4 in Section A.1.2 of the Appendix for details.
With RE and DB, we formulate the consistency of the initial workhorse Slasso estimator in the following lemma.
Lemma 2.
Under Assumptions 14, there exists some absolute constant $C_{{\rm m}}$ such that when the tuning parameter in (2.4) for the main regression satisfies $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$, as $n\to\infty$, we have
@@
Lemma 2 provides the consistency of the standardized coefficients, which is necessary to establish the asymptotic normality of XDlasso. The rates of tuning parameters are technical devices for proofs. In practice, we recommend cross-validation to select the tuning parameter $\lambda$.
@@ 3.2 The Auxiliary Regression
While Lemma 2 generalizes the convergence of Slasso from unit roots to LUR regressors, brand new theory must be developed for the auxiliary regression (2.12), again by the workhorse Slasso when $w_{j,t}$ is either LUR or stationary. Importantly, as we maintain an agnostic attitude about the persistence of the variable of interest, which is the dependent variable of the auxiliary regression, the theory must be applicable in a unified manner to accommodate two very different types of time series. Without loss of generality, let $\mathcal{M}_{x}=[p_{x}]$ and $\mathcal{M}_{z}=[p]\backslash[p_{x}]$ denote the integer sets indexing the locations of LURs and stationary regressors, respectively.
We first examine the case when $j\in\mathcal{M}_{x}.$ The standardized instrument $\widetilde{\zeta}_{j,t}$ has a different degree of persistence from $X_{-j,t}$ and $Z_{t}$, due to the different orders of integration. In the low dimensional framework, the FCLT in phillips2016robust yields the following rate of convergence in OLS:
@@
$\widetilde{\zeta}_{j,t}=X_{t}^{\top}0_{p_{x}}+Z_{-j,t}^{\top}\varphi_{z}^{(j)*}+\widetilde{\eta}_{j,t},$ (3.14)
where $\text{$\widetilde{\eta}$}_{j,t}=(\zeta_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$. Note that when $w_{j,t}$ is stationary, the IV $\zeta_{j,t}$ is close to $w_{j,t}$ under a large sample size, and thus asymptotically uncorrelated with the LUR regressors $X_{t}$. As a result, the coefficients associated with $X_{t}$ in the pseudo-true model (3.14) are zero.
In addition, the error term $\widetilde{\eta}_{j,t}$ is close to the stationary time series $(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$ and thus asymptotically uncorrelated to the nonstationary regressors $X_{t}$ due to different persistence. Furthermore, the coefficient $\varphi_{0z}^{(j)*}$ satisfies $\mathbb{E}\left(Z_{-j,t}(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})\right)=0$. Therefore, $\widetilde{\eta}_{j,t}$ is also asymptotically uncorrelated to the the stationary regressors $Z_{-j,t}$, thereby ensuring the consistency of the Slasso estimator $\widehat{\varphi}^{(j)}$ in (2.12). We impose the following assumption on the coefficient $\varphi_{0z}^{(j)*}$.
@@
$\|\varphi_{0z}^{(j)*}\|_{0}\leq s$ with $\varphi_{0z}^{(j)*}$ defined in (3.12) and $s$ specified in Assumption 4. Moreover, $\|\varphi_{0z}^{(j)*}\|_{1}\leq C_{1}$ for some absolute constant $C_{1}.$
To bound the LASSO estimation errors, we need sparsity of not only the main regression (2.1), but also the auxiliary regression (3.14). This is similar to the high dimensional sparse instrumental variable regression; see zhu2018sparse and gold2020inference. In Assumption 5, we slightly abuse the sparsity index $s$ to bound the number of nonzero coefficients in the vector $\varphi_{0z}^{(j)*}$. For simplicity, we directly impose this high-level sparsity assumption on $\varphi_{0z}^{(j)*}$, which can be deduced under the commonly used sparsity restriction on the precision matrix $\left(\mathbb{E}\left(Z_{t}Z_{t}^{\top}\right)\right)^{-1}$; see the definition of $s_{j}$ and the conditions in Theorem 2.1 of zhang2017simultaneous. Finally, the upper bound of the $L_{1}$-norm of $\varphi_{0z}^{(j)*}$ controls the variance of the error term in the pseudo-true model (3.14). Parallel conditions naturally hold for $j\in\mathcal{M}_{x}$ as the pseudo-true coefficients are zero in view of the local-to-zero OLS estimate displayed in (3.11).
The following proposition formally lays out the convergence rate of the auxiliary estimator (2.12).
@@
Suppose that Assumptions 15 hold. Then, there exists some absolute constant $C_{{\rm a},j}>0$ such that when $\mu_{j}=C_{{\rm a},j}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, we have
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right),$ (3.15)
@@
$\|n^{-1}\sum_{t=1}^{n}(D_{-j})^{-1}W_{-j,t}\widehat{r}_{j,t}\|_{\infty}\leq\dfrac{\mu_{j}}{2}\leq\dfrac{C_{{\rm a},j}(\log p)^{2+\frac{1}{2r}}}{2\sqrt{n^{(1-\tau)\wedge\tau}}},$ (3.16)
which is sufficient to bound the approximation error $(\widehat{{\rm B}}_{j}-{\rm B}_{j})$ in (2.7). Nevertheless, it is still necessary to establish the consistency of $\widehat{\varphi}^{(j)}$. This consistency result not only helps us show the asymptotic normality of the $t$-statistic $t_{j}^{{\rm XD}}$ defined in (2.15) to guarantee the asymptotic size of the test, but is also critical for the convergence rate of the standard error that governs the asymptotic power.
The consistency in the main equation and the auxiliary regression has paved the way for statistical inference. We will analyze the asymptotic size and power of the concerning test statistics in the next section.
3.3 Asymptotic Distributions
The desirable rate of the auxiliary regression built in Proposition 1 guarantees the following decomposition of the $t$-statistic:
$\frac{\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}}{\widehat{\omega}_{j}^{{\rm XD}}}={\rm sgn}_{j}\cdot\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}+O_{p}\left(\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\right),$ (3.17)
where ${\rm sgn}_{j}=\frac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is either 1 or 1-1 with probability one as $\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}$ is continuously distributed. The first term of the above expression is a counterpart of $\mathrm{N}_{j}/\omega_{j}$ in the discussion of generic desparsifying argument in (2.7). The second term is the convergence rate of the approximation error of the shrinkage bias analogous to $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}$. Under Assumption 4, the second term on the right-hand side of (3.17) is asymptotically negligible, thereby yielding the following asymptotic normality for any $j\in[p]$.
@@ 3.3 Asymptotic Distributions
Suppose Assumptions 1-5 hold. There exist absolute constants $C_{{\rm m}}$ and $C_{{\rm a},j}$ such that when $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\mu_{j}=C_{{\rm a},j}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, as $n\to\infty$ we have
@@
$\widehat{\omega}_{j}^{{\rm XD}}=\boldsymbol{1}\left\{j\in\mathcal{M}_{x}\right\}\cdot O_{p}(1/\sqrt{n^{1+\tau}})+\boldsymbol{1}\left\{j\in\mathcal{M}_{z}\right\}\cdot O_{p}\left(1/\sqrt{n}\right).$
The convergence rates displayed in Theorem 2 for the two types of regressors are coherent with the results in low dimensional predictive regressions. When $j\in\mathcal{M}_{x}$, the standard error converges faster than the rate $1/n^{\delta_{j}}$ for any $\delta_{j}\in(0,(1+\tau)/2)$. Thus, for the null hypothesis $\mathbb{H}_{0}:\beta_{j}^{*}=0$, the hypothesis testing based on the $t$-statistic $t_{j}^{{\rm XD}}$ is consistent under the alternative with $\beta_{j}^{*}=c/n^{\delta_{j}}$ for some $c>0$ over this class of $\delta_{j}$. The range $\delta_{j}\in(0,(1+\tau)/2)$ includes the important $1/\sqrt{n}$ rate of coefficients for the LUR regressor $X_{j,t}$, under which ${\rm var}(X_{j,t}\beta_{j}^{*})=O(1).$ The $1/\sqrt{n}$ factor thus balances the larger order of LUR regressors $X_{t}$ and the standard $O_{p}(1)$ stochastic order of $y_{t}$ (phillips2015halbert; lee2022lasso), and the test by XDlasso is consistent under this class of alternatives. When $j\in\mathcal{M}_{z}$, XDlasso achieves the standard $\sqrt{n}$-consistency.
@@
Unlike the convergence rates of Slasso in Lemma 2, the convergence rate of XDlasso in Theorem 2 is independent of the variable dimension $p$ and the sparsity index $s$. In view of (2.7), in XDlasso the order of the approximation error of the shrinkage bias $(\widehat{{\rm B}}_{j}-{\rm B}_{j})$ depends on $p$ and $s$, but it is dominated by the order of the noise component ${\rm N}_{j}$. The convergence rate of XDlasso therefore follows the order of ${\rm N}_{j}$, which relates only to the sample size $n$.
Finally, when we are interested in a joint null hypothesis $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}}$ for the coefficients indexed by $\mathcal{J}\subset[p]$ with a fixed cardinality, we can use the Wald statistic in (2.19) for this test.
Theorem 3.
Suppose that $|\mathcal{J}|$ is fixed as $n\to\infty$, and the conditions in Theorem 1 hold for all $j\in\mathcal{J}$. Under the null hypothesis $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}}$, we have
${\rm Wald}_{\mathcal{J}}^{{\rm XD}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\chi_{|\mathcal{J}|}^{2}.$
Theorem 3 is a natural extension from the $t$-statistic for a one-dimensional univariate hypothesis to a simultaneous multivariate hypothesis. Notice that the Wald statistic is valid even if $\mathcal{J}$ includes a mixture of LUR regressors and stationary ones. It allows the researcher to maintain the agnostic attitude when conducting the joint hypothesis testing.
Let us summarize the insights gained from the theoretical development. The fundamental principle behind the Dlasso method, as discussed in zhang2014confidence, is based on the Frisch-Waugh-Lovell theorem. This theorem purges the influence of other control variables to overcome shrinkage bias. In the context of persistent predictors, the two-stage least squares approach, as developed in magdalinos2009limit is used to eliminate Stambaugh bias. Each piece of XDlasso is a machine learning version of a classical idea. To adapt these existing procedures to high dimensional predictive regressions, we must rely on the consistency of Slasso, with its extension to LURs. This consistency is crucial for both the main regression and the auxiliary regression. Moreover, as mentioned in Remark 8, the construction of the IV and the choice of $\tau$ must be tailored to balance the two types of predictors. This approach is original in that such a restriction is unique to high dimensional models and has not been studied before even in conventional low dimensional settings.
4 Monte Carlo Simulations
In this section, we evaluate the performance of the proposed XDlasso inference procedure by comparing its test size and power with those of Dlasso. Although Dlasso and other existing LASSO bias-correction procedures are not designed to handle persistent regressors, we include this comparison to highlight the value added by the IVX transformation. Furthermore, to demonstrate the robustness of the XDlasso approach, we benchmark it using tests based on infeasible estimators, where an oracle reveals the locations of the nonzero coefficients.